A is 102 years younger than r. 4yrs ago r was 4x as age. How old is a now.I
Answer by ankor@dixie-net.com(22740)

Kind of ridiculous but we can do it
:
A is 102 years younger than r.
A = r - 102
4yrs ago r was 4x as age.
r - 4 = 4(A - 4)
r - 4 = 4A - 16
r = 4A - 16 + 4
r = 4A - 12
from the 1st equation, replace A with (r-102)
r = 4(r - 102) - 12
r = 4r - 408 - 12
r = 4r - 420
420 = 4r - r
420 = 3r
r = 420/3
r = 140 yrs old
:
How old is a now.
A = 140 - 102
A = 38 yrs old
:
:
See if that checks out for the statement
"4 yrs ago r was 4x A's age. "
136 = 4(34)